# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

### v1

@@ 5 Experiments
In this section, we exemplify the benefits and variability of the proposed acquisition strategy by performing experiments across three categories that differ substantially from one another. These problems originate from distinct literature within BO and BED, and none of them are canonical BO or BED tasks. Thus each of them is evaluated against a different set of BO-type (optimization-focused) and BED-type (learning-focused) baselines appropriate for that specific task, ensuring a fair and rigorous comparison.
Figure 1: Performance evaluation for constrained system identification on environmental monitoring in 2d plume fields. Error bars represent $\pm 1$ std over 5 seeds.
5.1 Parametric Models with Known Invariant Goals
We first consider a simplest task where the goals are known and invariant, and we have a decent understanding of the model, where only certain finite parameters are unknown.
One typical problem structure is the constrained system identification whose objective is to precisely learn unknown system parameters, $\theta$, when valid observations can only be gathered under specific operational constraints, quantified by $C(y)\leq 0$. For instance, when monitoring a chemical plume, the goal is to identify its source ($\theta$) while ensuring that sensor measurements ($y$) do not exceed a saturation threshold ($C(y)$).
In such cases, it is straightforward to choose the interested state as $s=\theta$, and the energy function as $h(y|\mathcal{D}_{t})=\mathbb{I}(C(y)>0)$. This leads to an acquisition function:
$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(\theta;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}\left[\mathbb{I}(C(y)>0)\right].\end{split}$
Tasks. We perform experiments on a real-world environmental monitoring problem in 2d plume fields where the sensors have a saturation threshold $y_{max}$ (i.e., $C(y)=y-y_{max}$) (Detailed settings and hyper-parameter choices in Appendix D.2). We consider three types of monitoring tasks: (a) locating the unknown source location; (b) estimating unknown wind direction and strength; and (c) identifying the active sources in fields with multiple sources.
Baselines. We compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by choosing the point that leads to the least probability of violating the constraint (BO-type); and (c) EIG about the unknown parameter (BED-type).
Evaluations. We evaluate the performance from both epistemic and pragmatic perspectives: (a) estimation accuracy; and (b) constraint violations.
Results. Fig. 1 shows that our approach achieves consistently stronger query efficiency over baselines while respecting all operational constraints, with cumulative constraint violation always being zero. This advantage is especially evident in the source localization task, where the drive for information and the need to satisfy constraints create opposing pressures. Resolving this conflict, our approach reaches near-perfect estimation using up to 40% fewer queries than competing methods.
@@ 5.1 Parametric Models with Known Invariant Goals
Figure 2: Performance evaluation for targeted active search on failure discovery in autonomous driving scenarios. Error bars represent $\pm 1$ std over 4 seeds.
5.2 Non-Parametric Models with Known Evolving Goals
Next, we consider a more challenging setting where the task condition evolves, and the model is fully black-box such that we need to resort to a non-parametric model (e.g., GPs).
One such example is the targeted active search in multi-objective design problems. The objectives are viewed as metrics where specific ranges carry particular significance, and the goal is to design experiments that maximize coverage of these important regions $\mathcal{S}$.
Adopted from Malkomes et al. (2021), we assume that the experimental design problems possess a sense of known resolution, such as simulation accuracy or manufacturing precision/tolerance, and any outcome within distance $\delta$ of another does not convey extra information about $\mathcal{S}$. Similar to Malkomes et al. (2021), we define the coverage neighborhood of any $y$ as $\mathbb{C}_{\delta}(y):=\{y^{\prime}:d(y,y^{\prime})<\delta\}$. And the coverage neighborhood of a set of points $Y$ is defined as $\mathbb{C}_{\delta}(Y):=\bigcup_{y\in Y}\mathbb{C}_{\delta}(y)$.
Since the goal is to cover as much volume of the interested region $\mathcal{S}$ as possible, $h(y|\mathcal{D}_{t})$ can be chosen as $h(y|\mathcal{D}_{t})=\text{Vol}(\mathcal{S})-\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})$, where $Y$ contains all the history of outcomes. Similarly, coverage over the design configurations can be flexibly incorporated by expanding the interested state into the whole input space, i.e., $s=f_{\mathcal{X}}$. With non-parametric models like GPs, we can leverage Lemma 3.2 and derive an acquisition function:
@@ 5.2 Non-Parametric Models with Known Evolving Goals
$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(f_{x};y|\mathcal{D}_{t})+\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})].\end{split}$
Tasks. We perform experiments on a real-world failure discovery problem in autonomous driving scenarios where the perception module, a YOLO detector, may fail due to multiple causes, which could potentially lead to collisions (3d input-2d output). We consider three target sets with decreasing volume, i.e., Target Set 1 $\supset$ Target Set 2 $\supset$ Target Set 3 (Detailed settings and hyper-parameter choices in Appendix D.3).
Baselines. We again compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by maximizing coverage volume in metrics space (BO-type); and (c) EIG by maximizing coverage volume in parameter space (BED-type).
@@
Evaluations. We evaluate the performance by inspecting the coverage volume over both spaces: (a) metrics space ($\mathbb{C}_{\delta}(Y)$); and (b) parameter space ($\mathbb{C}_{\delta}(X)$).
Results. As shown in Fig. 2, our AIF algorithm effectively balances coverage of both the parameter and metrics spaces. This capability is especially impactful for smaller target sets where the search is more difficult. In the most challenging case (Target Set 3), our method identifies nearly 10% more of the critical failure region compared to the leading baseline.
@@
Figure 3: Performance evaluation for composite BO with unknown preferences. Error bars represent $\pm 1$ std over 20 seeds for vehicle safety, penicillin, and 5 for energy resource allocation.
5.3 Non-Parametric Models with Unknown Goals
Finally, we look into the most difficult setting where both the models and goals are black-box.
One practical scenario arises from the composite Bayesian optimization in multi-objective optimization problems. The objectives are weighted by a preference function $g(y)$ that is unknown a priori and must be simultaneously learned during the optimization process.
As assumed in Lin et al. (2022), we can estimate the unknown preference function $g(y)$ by asking a decision-maker (DM) to express preferences over pairs of outcomes $(y_{1},y_{2})$. Let $z(y_{1},y_{2})\in\{1,2\}$ indicate whether the DM preferred the first or second outcome offered. Following Chu and Ghahramani (2005), we assume that the DM’s responses are distributed according to a probit likelihood: $L(z(y_{1},y_{2})=1|g(y_{1}),g(y_{2}))=\Phi(\frac{g(y_{1})-g(y_{2})}{\sqrt{2}\lambda}),$ where $\lambda$ is a hyperparameter, and $\Phi$ is the standard normal CDF.
Taking the joint exploration and exploitation of the preference function $g(y)$ into consideration, we can choose $h(y|\mathcal{D}_{t})$ as the EFE of the preference model, i.e., $h(y|\mathcal{D}_{t})=\text{EFE}(g(y))$, which naturally extends the acquisition function into a nested structure:
[…]
where $x=[x_{1},x_{2}]$ are two jointly evaluated candidates, and $\gamma,\beta\geq 0$.
Tasks. We perform experiments on three real-world problems, including vehicle safety (5d-3d), penicillin production simulator (7d-3d), and distributed energy resource allocation in power grids (40d-4d) (Detailed settings and hyper-parameter choices in Appendix D.4).
Baselines. Different from BOPE in Lin et al. (2022), which separates and iterates the preference exploration and experimentation stages, our approach unifies the exploration and exploitation of both outcome and preference models simultaneously. In its best setting (i.e., switch stages and update models at each step), their acquisition strategy reduces to a special case of our proposed (4) with $\beta=\gamma=0$. Thus, we design a set of baselines to investigate the effect that each component in (4) plays for the joint exploration and exploitation of hierarchical models. We ablate the full AIF strategy (g + g IG + f IG) by (a) removing the last term (EIG); (b) removing the first term (g + g IG); and (c) removing the first two terms (g only).
@@ 5.3 Non-Parametric Models with Unknown Goals
Evaluations. We evaluate their performance by inspecting the best preference of all collected outcomes using the true preference function $g(y)$.
Results. Fig. 3 illustrates our AIF approach’s outstanding ability to learn unknown preference functions, a key advantage over baselines that often fail due to ill-directed queries. As the tasks become more complex and noisy (from (a) to (c)), our results demonstrate that every component of our acquisition function (4) plays an irreplaceable role in achieving optimal performance. This benefit is most stark in the energy resource allocation task, where competing methods failed to capture any meaningful preference model while our approach consistently succeeded.
Benefits of joint learning and optimization. To highlight the benefits of jointly learning and optimizing, rather than separating these into stages, we compare our method against several BOPE variants from Lin et al. (2022) that use different stage-wise design choices. Detailed experimental setups, design choices, plots, and analyses are provided in Appendix D.4.3. The results show that our method naturally balances exploration and exploitation at each step and consistently discovers higher-preference regions, whereas the BOPE variants are highly sensitive to how the stages are configured. Consequently, stage-wise approaches like BOPE require careful manual tuning of these choices, while our unified formulation automates this trade-off and is therefore more amenable to higher-order hierarchical models.
6 Conclusion and Limitation
We presented pragmatic curiosity, an AIF-based paradigm for hybrid learning-optimization under expensive black-box evaluations. By selecting actions via EFE minimization, the approach unifies goal-seeking and information-seeking within a single acquisition objective. Across constrained system identification, targeted active search, and composite optimization with unknown preferences, pragmatic curiosity consistently outperformed strong BO-type and BED-type baselines, improving estimation accuracy, critical-region coverage, and final solution quality under fixed budgets.
The limitation of our approach stems from the specification of the problem-dependent energy/preference model; misspecification can bias acquisition and degrade performance. Performance also inherits assumptions of the underlying surrogate and observation models; strong model mismatch or non-stationarity can compromise both uncertainty quantification and preference-induced guidance. Future work includes extending the evaluation of this paradigm to multi-agent, long-horizon, or multi-fidelity settings.
Impact Statement
This work advances sample-efficient decision-making for engineering and scientific systems where evaluations are expensive and objectives are coupled to uncertainty. By unifying goal-seeking and information-seeking within a single EFE objective, pragmatic curiosity provides a general acquisition-design template that can reduce experimental cost, accelerate discovery, and improve reliability in workflows such as safe system identification, targeted failure-region discovery, and preference-driven design. In domains like environmental monitoring, autonomous-system testing, and energy-resource allocation, the practical benefit is fewer costly or risky evaluations for a given level of performance and confidence, which can translate into reduced resource consumption and safer experimentation.
Potential negative impacts stem from misuse or mis-specification of the preference/energy model: poorly chosen goals or biased preferences can steer exploration toward harmful regions (e.g., unsafe operating conditions) or encode inequitable outcomes in preference-driven optimization. In addition, the ability to more efficiently identify system vulnerabilities could be misused in adversarial settings. These risks motivate careful deployment practices, including explicit safety constraints, transparent preference specification, robustness checks under model mismatch, and domain-appropriate oversight when the method is applied to safety-critical or socially sensitive applications.
Appendix A Related Work
Several independent attempts have been made to handle the hybrid objectives of goal seeking and learning. Below, we discuss some key prior works, originating from three viewpoints: (1) BO-based settings, (2) BED (also known as Bayesian active learning, BAL) based works, and (3) some works that have acknowledged the synergy and attempted to present a unifying viewpoint, using adaptive sampling and AIF-based approaches.
On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems by minimizing the ratio between squared expected single-period regret and a measure of information gain. Hvarfner et al. (2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the Gaussian process hyperparameter even as it searches for the optimum. Tu et al. (2022); Hvarfner et al. (2022a) propose an information-theoretic acquisition function for multi-objective BO, which considers the joint information gain for the optimal set of inputs and outputs. Overall, these works acknowledge and demonstrate the value of information gain in enhancing the capabilities of goal-seeking strategies, with an ultimate focus on optimization. However, they do not explicitly comment on the synergy of goal seeking and learning, or formalize the acquisition strategies as a way to explicitly balance both objectives.
On the BED (also known as Bayesian active learning, BAL) side, Smith et al. (2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution. Lookman et al. (2019) discusses the performance of various BAL and BO strategies for material design application, noting their connection with adaptive sampling, and simultaneous exploration and exploitation. This shows that the realistic application of BO/BED-based approaches often necessitates a formalized viewpoint unifying each distinct objective.
These efforts begin to show the growing synergy between BO and BED techniques, where information-gain criteria are leveraged to enhance optimization and vice versa. In search of a unifying perspective for these techniques, Di Fiore et al. (2023) views BO and BAL as symbiotic adaptive sampling techniques driven by common principles of ‘goal-driven learning’. However, their viewpoint is limited to BO and BED as separate realizations of adaptive sampling, and they do not discuss a way to combine both objectives in a single acquisition strategy.
In this paper, we present a unified perspective that arises naturally from the principles of Active Inference (AIF) (Friston, 2010; Friston et al., 2017). Originally developed in computational neuroscience, AIF has been successfully applied in fields such as robotics (Lanillos et al., 2021), cognitive science (Smith et al., 2022), ecosystem modeling (Friston et al., 2024), and goal-directed planning in partially observable environments (Matsumoto et al., 2025). These works leverage the ability of AIF to accommodate goal-directed and epistemic considerations, which we utilize as the backbone for our unifying mechanism discussed in this paper.
Appendix B Proof of Theorem 3.1
Proof.
Recall that
$G=\underbrace{-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]}_{\text{Term 1}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y|x)}_{\text{Term 2}}$
Since our purpose is to align the true outcome distribution given the decision $x$ (i.e., $p(y|x)$) with the prior preference distribution $p(y)$, we replace the second term of $G$ by $\mathbb{E}_{q(y|x)}\log p(y)$. It is commonly referred to as pragmatic value, which scores the anticipated surprisal of the future outcomes given a predictive distribution $q(y|x)$.
For the first term of $G$, add and subtract $\log q(s|x,y)$ inside the expectation, we get
$\begin{split}\text{Term 1}&=-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)\right]+\mathbb{E}_{q(s|x,y)}\left[\log q(s|x,y)-\log q(s|x)\right]\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\underbrace{-D_{\text{KL}}(q(s|x,y)||p(s|x,y))}_{\text{Term 3}}+\underbrace{D_{\text{KL}}(q(s|x,y)||q(s|x))}_{\text{Term 4}}\right].\end{split}$ (5)
Since the surrogate model $q(s)$ usually will not be updated until a new outcome $y$ is observed, it doesn’t depend on the decision variable $x$ solely, i.e., $q(s|x)=q(s)$. Thus Term 4 becomes $D_{\text{KL}}(q(s|x,y)||q(s))$, which is commonly referred to as the epistemic value, or the expected information gain of a state when it is conditioned on expected observations.
Term 3 is, in general, intractable since it depends on the true posterior distribution $p(s|x,y)$ that is often unknown. The best surrogate model we can have is $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is constructed from all available data $\mathcal{D}_{t}$. Thus if we treat this surrogate model $q(s)$ as the best approximation of the true model $p(s)$, i.e., $q(s)\approx p(s)$, and the surrogate and true models both follow the Bayes’ rule, i.e., $q(s|x,y)=\frac{p(x,y|s)q(s)}{\int p(x,y|s)q(s)ds},\quad p(s|x,y)=\frac{p(x,y|s)p(s)}{\int p(x,y|s)p(s)ds},$ then Term 3 vanishes. ∎
Appendix C Proof of Lemma 3.2

### v2

@@ 5 Experiments
PraC provides a principled but flexible acquisition-design template through three coupled choices: a symbolic representation of what is worth learning, a regret-based potential that encodes what outcomes matter, and a curiosity coefficient that determines how strongly epistemic and pragmatic pressures are traded against one another. In this section, we instantiate these choices across three regimes of increasing complexity.
The experiments are organized according to the structure exposed by PraC. We begin with fixed global symbols and known goals, where the latent quantity is a finite hypothesis and the regret is a downstream Bayes risk. We then consider induced local symbols with known but evolving goals, where the relevant latent quantity is induced by a non-parametric surrogate and the goal depends on the coverage already obtained. Finally, we study a hierarchical setting with unknown goals, where the agent must learn both how actions produce outcomes and how those outcomes should be valued. We study dynamic scheduling of $\beta_{t}$ in detail in the first regime, where its closed-loop interpretation is most transparent. In the non-parametric and hierarchical regimes, we use fixed curiosity weights to isolate the effects of representation and regret design.
Figure 1: Decision-oriented plume-monitoring performance over 20 seeds. Top row: posterior Bayes risk. Bottom row: task-specific realized decision metric: response loss for source response localization and consequence-weighted dispatch, and missed-source risk for active source prioritization. Shaded regions show mean $\pm$ standard error.
5.1 Fixed Global Symbols with Known Goals
We begin with the most transparent regime, in which the symbolic structure is fixed and the downstream goal is explicitly specified. Here the unknown quantities are finite-dimensional hypotheses with stable semantic meaning across all decisions, so the epistemic target is a fixed global symbol. At the same time, the task objective is known in advance and can be expressed through a closed-form downstream regret. This regime provides the cleanest setting in which to examine whether curiosity can be elevated from a heuristic coefficient to a monitored and interpretable part of the decision process.
A representative problem of this form is decision-oriented monitoring, where the purpose of sensing is not merely to estimate environmental parameters, but to improve a downstream response decision. At each iteration, the agent maintains a posterior over latent environmental hypotheses $\theta$, selects a sensor query $x$, observes a plume count $y$, updates the posterior, and evaluates the downstream decision induced by the updated belief. Let $a\in\mathcal{A}$ denote a downstream action, and let $L(a,\theta)$ denote the task-specific loss incurred by taking action $a$ when the true hypothesis is $\theta$. Given the current surrogate posterior $q_{t}(\cdot)$, the Bayes risk is
$\mathrm{BR}(\mathcal{D}_{t})\coloneq\min_{a\in\mathcal{A}}\mathbb{E}_{\theta\sim q_{t}}\left[L(a,\theta)\right].$
This measures the downstream loss the agent still expects to incur if it acts optimally under its current belief.
For a candidate sensor query $x$, define the expected Bayes-risk reduction
$\widehat{\Delta\mathrm{BR}}_{t}(x)\coloneq\mathrm{BR}(\mathcal{D}_{t})-\mathbb{E}_{q_{t}(y\mid x)}\left[\mathrm{BR}(\mathcal{D}_{t}\cup\{(x,y)\})\right].$
The PraC acquisition becomes
$\alpha_{t}(x)=\beta_{t}I(\theta;(x,y)\mid\mathcal{D}_{t})+\widehat{\Delta\mathrm{BR}}_{t}(x).$ (4)
Tasks. We evaluate this setting on three decision-oriented plume-monitoring tasks in 2D fields. (a) Source response localization asks the agent to localize a plume source well enough to dispatch a response team near it. Here the downstream action $a$ is a response location, and the loss $L(a,\theta)$ is the normalized squared distance between the response location and the source location specified by $\theta$. (b) Consequence-weighted dispatch uses the same response-decision structure, but errors in high-consequence regions are assigned larger loss. (c) Active source prioritization asks the agent to select which suspected industrial leaks to repair under a limited crew budget. Here the downstream action $a$ is a subset of sources, and the loss $L(a,\theta)$ is the weighted risk of true active sources not selected for repair. Detailed plume models, hypothesis spaces, downstream losses, and Bayes-risk computations are provided in Appendix C.2.
Baselines. We compare two variants of PraC with (i) dynamic $\beta$ as scheduled in Section 4.3, and (ii) fixed-$\beta$ with $\beta=5$, against (a) random sampling, (b) pure information gain (EIG), and (c) decision-greedy policy by minimizing $BR_{t}(x)$.
Evaluations. We report posterior Bayes risk and task-specific realized decision metrics. For source response localization and consequence-weighted dispatch, the realized metric is response loss under the true source. For active source prioritization, the realized metric is missed-source risk, namely the total consequence weight of true active sources not selected for repair. Additional metrics, including response success, parameter error, top-$k$ recall, weighted top-$k$ recall, and detailed analysis are reported in Appendix C.2.6.
Results. Fig. 1 shows that dynamic PraC gives the strongest overall performance across the three decision-oriented plume tasks. In source response localization, dynamic PraC achieves the lowest final Bayes risk while matching the best realized response loss. In consequence-weighted dispatch, decision-greedy obtains the lowest final Bayes risk, but dynamic PraC obtains the best realized response loss and the lowest parameter error. This distinction is important: Bayes risk measures expected loss under the agent’s current posterior, whereas realized response loss evaluates the downstream decision under the true latent environment. Dynamic curiosity can therefore sacrifice some immediate posterior-risk reduction in order to collect information that improves the realized downstream decision. In active source prioritization, dynamic PraC obtains the lowest final Bayes risk, improving over fixed-$\beta$ PraC, decision-greedy, EIG, and random sampling. Dynamic and fixed-$\beta$ PraC tie on missed-source risk, while both substantially improve over EIG and random sampling.
Dynamic scheduling of curiosity. Scheduler diagnostics are reported in Appendix C.2.7. The feedforward scale estimates the local exchange rate between Bayes-risk reduction and information gain, while the feedback activation suppresses curiosity as decision-relevant uncertainty decreases. These diagnostics support the intended interpretation of $\beta_{t}$: dynamic PraC does not merely reduce environmental uncertainty for its own sake, but learns distinctions that matter for downstream response decisions.
@@ 5.1 Fixed Global Symbols with Known Goals
Figure 2: Performance evaluation for targeted active search on failure discovery in autonomous driving scenarios. Top row: coverage in the outcome/metric space. Bottom row: coverage in the input/parameter space. Error bars represent $\pm 1$ standard deviation over 4 seeds.
5.2 Induced Local Symbols with Known Evolving Goals
We now move to a more challenging regime in which the latent structure is no longer naturally captured by a fixed finite-dimensional parameter. Instead, the relevant symbolic representation is induced locally by the contemplated observations, as described in Section 4.1. At the same time, the goal remains known, but it is no longer invariant: it evolves with the outcomes already collected. This regime is designed to highlight the value of PraC’s symbolic flexibility. To isolate that effect, we retain a fixed curiosity weight in this subsection.
A representative problem of this form is targeted active search in multi-objective design. Here the objective is not simply to optimize a scalar reward, but to cover as much as possible of an outcome region $\mathcal{S}$ deemed important. Following Malkomes et al. (2021b), we assume a known resolution $\delta$, so that outcomes within distance $\delta$ are treated as redundant with respect to target coverage. Define the coverage neighborhood of an outcome $y$ and a set of outcomes $Y$ by
$\mathbb{C}_{\delta}(y):=\{y^{\prime}:d(y,y^{\prime})<\delta\},\qquad\mathbb{C}_{\delta}(Y):=\bigcup_{y\in Y}\mathbb{C}_{\delta}(y).$
Then a natural potential is
$h_{t}(y)=\mathrm{Vol}(\mathcal{S})-\mathrm{Vol}\!\left(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S}\right),$
which penalizes outcomes that add little new coverage of the target set. Meanwhile, the relevant epistemic target is the local non-parametric symbol induced by the contemplated observation, so we take $s=f_{\mathcal{X}}$ and use the representation-compression result in Lemma 4.1 to obtain
@@ 5.2 Induced Local Symbols with Known Evolving Goals
$\alpha(x\mid\mathcal{D}_{t})=\beta I(f_{x};y\mid\mathcal{D}_{t})+\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}\!\left[\mathrm{Vol}\!\left(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S}\right)\right].$
Tasks. We study failure discovery for a YOLO-based perception module in autonomous-driving scenarios. The input is a 3-dimensional scenario parameterization, and the outcome is a 2-dimensional failure metric describing two failure modes that may lead to collision. We consider three nested target sets of decreasing volume, $\mathcal{S}_{1}\supset\mathcal{S}_{2}\supset\mathcal{S}_{3},$ where smaller target sets correspond to rarer and more difficult failure regions. Details of the CARLA simulation, failure metrics, and target-set definitions are provided in Appendix C.3.
Baselines. We compare PraC against random sampling and two expected coverage improvement (ECI) variants adapted from Malkomes et al. (2021b): (a) ECI-y, which greedily prioritizes target coverage in the outcome/metric space, and (b) ECI-x, which prioritizes coverage in the input/parameter space. These baselines isolate the two coverage pressures that PraC jointly coordinates.
@@
Evaluations. We evaluate coverage in both spaces, including metric-space coverage $\mathbb{C}_{\delta}(Y)$, and parameter-space coverage $\mathbb{C}_{\delta}(X)$.
Results. As shown in Fig. 2, PraC effectively balances the two forms of coverage. It does not merely spread samples broadly in parameter space, nor does it myopically chase already discovered target regions in outcome space. Instead, the information term encourages learning of the local surrogate structure, while the coverage potential directs samples toward still-uncovered parts of the target region. This coordination is especially useful for smaller target sets, where informative failures are sparse and purely greedy coverage strategies become brittle. In the most challenging case, Target Set 3, PraC discovers nearly 10 percentage points more normalized target-region coverage than the strongest baseline.
Role of induced symbols. This experiment emphasizes the representational role of PraC. Once the epistemic target is allowed to be induced locally by the contemplated observation, the same acquisition principle naturally adapts from finite-hypothesis parameter identification to black-box target search with non-parametric surrogates.
@@
Figure 3: Performance evaluation for composite BO with unknown preferences. Metric is the best preference value attained among all collected outcomes. Error bars represent $\pm 1$ standard deviation over 20 seeds.
Figure 4: Comparison with BOPE-style stage-wise variants for energy resource allocation. Metric is the best preference value attained. Error bars represent $\pm 1$ standard deviation over 20 seeds.
5.3 Hierarchical Symbols with Unknown Goals
Finally, we consider the most difficult regime, in which neither the predictive structure nor the goal is fully specified in advance. In this setting, the agent must learn not only how actions produce outcomes, but also how outcomes should be valued. This experiment is designed to emphasize the role of regret in PraC: the pragmatic term is no longer specified directly, but is itself inferred through an additional layer of modeling. To keep the focus on this hierarchical regret design, we retain fixed curiosity weights in this subsection.
A representative example is composite Bayesian optimization with unknown preferences. The outcomes are multi-objective vectors, but the scalar preference over them is determined by an unknown function $g(y)$ that must be learned during optimization. Following Lin et al. (2022), we infer $g(y)$ from pairwise preference queries. For a pair of outcomes $\tilde{y}=[y_{1},y_{2}]$, let $l(y_{1},y_{2})\in\{1,2\}$ denote which outcome is preferred by the decision-maker. Following Chu and Ghahramani (2005), we assume the probit likelihood
$\Pr(l(y_{1},y_{2})=1\mid g(y_{1}),g(y_{2}))=\Phi\!\left(\frac{g(y_{2})-g(y_{1})}{\sqrt{2}\lambda}\right),$
where $\lambda$ is a hyperparameter and $\Phi$ is the standard normal CDF. Under this formation, outcome $y_{1}$ is more likely to be preferred when it has lower latent regret than $y_{2}$.
At each iteration, the method selects a candidate pair $\tilde{x}=[x_{1},x_{2}]$ and observes the corresponding pair of outcomes $\tilde{y}=[y_{1},y_{2}]$ and a preference label $l$. This yields the hierarchical PraC acquisition
$\begin{split}\alpha(\tilde{x}\mid\mathcal{D}_{t})=\beta_{t}I(f_{\mathcal{X}};(\tilde{x},\tilde{y})\mid\mathcal{D}_{t})+\mathbb{E}_{q(\tilde{y}\mid\tilde{x},\mathcal{D}_{t})}\left[\gamma_{t}I(g_{\mathcal{Y}};(\tilde{y},l)\mid\mathcal{D}_{t})-\min_{y\in\{y_{1},y_{2}\}}\mathbb{E}_{q(g_{y}\mid y,\mathcal{D}_{t})}[g_{y}]\right],\end{split}$ (5)
where the first term learns the outcome model $f$; the second term learns the regret model $g$ through informative pairwise comparisons; and the final term exploits the currently predicted best member of the evaluated pair by favoring low expected regret.
Tasks. We evaluate this setting on three real-world multi-objective optimization problems of increasing complexity: (a) vehicle safety with 5-dimensional inputs and 3-dimensional outcomes, (b) penicillin production with 7-dimensional inputs and 3-dimensional outcomes, and (c) energy allocation in power grids with 40-dimensional inputs and 4-dimensional outcomes. Details of the simulators, outcome definitions, and ground-truth preference models are provided in Appendix C.4.
Baselines. We first compare the full hierarchical PraC acquisition against ablated variants obtained by removing: (a) the outer epistemic term over $f$, (b) the inner epistemic term over $g$, or (c) both epistemic terms, leaving only the pragmatic preference term. These ablations isolate the contribution of outcome-model learning, preference-model learning, and their joint interaction. We additionally compare against BOPE-style stage-wise variants (Lin et al., 2022) for the energy-allocation task, where the distinction between joint and staged learning is most pronounced. Detailed design of BOPE-style variants are provided in Appendix C.4.3.
@@ 5.3 Hierarchical Symbols with Unknown Goals
Evaluations. We evaluate the best preference value attained among all collected outcomes using the true latent preference function.
Results. Fig. 3 shows that PraC consistently outperforms the ablated variants in learning and exploiting the unknown preference structure. As the tasks become more complex and noisier, each component of the hierarchical acquisition becomes increasingly important. This is most evident in the energy-allocation problem, where variants that remove either outcome-model exploration or preference-model exploration fail to discover high-preference regions reliably, whereas the full PraC acquisition continues to improve.
Benefits of joint learning and optimization. Fig. 4 compares PraC with BOPE-style stage-wise variants on energy allocation. BOPE alternates between preference exploration and experimentation through manually specified stages or switching rules. PraC instead integrates outcome learning, preference learning, and preference-guided optimization at every iteration through a single acquisition. The comparison shows that stage-wise methods are sensitive to their switching schedules, while PraC more reliably discovers high-preference regions. Thus, the hierarchical version of PraC not only learns what the world does, but also learns what matters, without requiring a manually staged separation between learning and optimization.
6 Conclusion
This paper introduced Pragmatic Curiosity (PraC), an active-inference framework for hybrid learning and optimization. PraC addresses settings in which an agent cannot cleanly separate what it needs to learn from what it needs to do. By coupling epistemic information-seeking with a regret-based pragmatic potential, PraC provides a unified acquisition principle for deciding which information is worth acquiring because of how it may improve action.
The framework exposes three operational design choices: symbols, the latent quantities whose clarification can change action; regrets, the potential landscapes that encode task value or shortfall; and curiosity, the exchange rate between information and pragmatic value. These choices provide a common language for interpreting existing BO and BED acquisition rules, while enabling new designs for hybrid problems where learning, optimization, and preference formation are coupled.
We instantiated PraC across decision-oriented monitoring, targeted active search, and composite Bayesian optimization with unknown preferences. Across these regimes, PraC reduced downstream decision risk, improved coverage of critical outcome regions, and jointly learned predictive and preference structures without relying on task-specific staging rules.
More broadly, PraC suggests that curiosity should not be treated merely as an exploration heuristic. When grounded in task-relevant symbols and regulated by regret-based potentials, curiosity becomes a mechanism for refining knowledge and action together. Future work includes non-myopic extensions, stronger theoretical guarantees, scalable approximations, and applications to embodied robotic systems.
Appendix A Proof of Lemma 4.1 (Representation Compression)